OpenStudy (callisto):

Statistics question #2

OpenStudy (callisto):

How should I start?

OpenStudy (anonymous):

the numbers are probably in ascending order

OpenStudy (callisto):

That's the thing I'm not sure of...

OpenStudy (anonymous):

I mean ... the expressions should be ordered

OpenStudy (anonymous):

ascending or descending

OpenStudy (anonymous):

in this case x = 11

OpenStudy (callisto):

But it should be like this: x+17> x-1 > x-3 > x-4 ?! Oh... seems I got something from here..

OpenStudy (callisto):

Alright, I got it

OpenStudy (dumbcow):

:)

OpenStudy (unklerhaukus):

\[\langle x\rangle=\sum\limits_{n=0}^{\infty} xP(x)=\sum\limits_{n=0}^{\infty} x\frac{N(x)}{N}\] \[8=\frac{1}{5}\sum\limits_{n=0}^{6} xN(x)\]\[8=\frac{1}{5} \left((15\times1)+(x-1) \times 1+(x-3)\times1 +(x-4)\times1+(x-17)\times 1 \right)\]

OpenStudy (thomas5267):

The mean is 19. Isn't it?

OpenStudy (callisto):

@thomas5267 Nope. That's not in the option

OpenStudy (dumbcow):

should be 12

OpenStudy (callisto):

x+17> x-1 > x-3 > x-4 15 could be the largest or the smallest Case I: if 15 is the largest 15>x+17> x-1 > x-3 > x-4 Median = x-1 =8 => x=9 mean = 12 Case II: if 15 is the smallest x+17> x-1 > x-3 > x-4 >15 median = x-3 = 8 => x=11 mean = 13.6 .... Oh fail.... I thought I got 12 as the only answer.. @dumbcow Time to explain~

OpenStudy (mani_jha):

If I assume x to be a positive integer, the descending order would be: x+17,15,x-1,x-3,x-4

OpenStudy (dumbcow):

15 can't be smallest, if x=11 then x-4 < 15

OpenStudy (dumbcow):

actually 15 can't be largest either, if x=9 then x+17 > 15 order should be x+17 > 15> x-1 > x-3 > x-4

OpenStudy (thomas5267):

So that we are assuming the question only gives us positive integer. @dumbcow

OpenStudy (dumbcow):

no we are not assuming that but given the median is 8, it is impossible for x to be negative x can only be 9

OpenStudy (anonymous):

found this : Mode = Mean—3 (Mean—Median) the mode in your sample is 0

OpenStudy (anonymous):

0 =M -3 (M-8)

OpenStudy (dumbcow):

never seen that before...will that work in the general sense? what if mean = median, there is no guarentee mode = mean in that case

OpenStudy (anonymous):

I haven't seen this also ... but I hated Stats.

OpenStudy (anonymous):

does 12 work ?

OpenStudy (dumbcow):

1,2,3,4,5 mean = 3 median = 3 mode = 0 formula breaks down ? :(

OpenStudy (callisto):

12 is the answer

OpenStudy (anonymous):

In a symmetrical distribution, mean, median and mode are identical and have the same value. However, in actual life most distributions are not symmetrical—they are skew. (These concepts of symmetry and skewness would be discussed in the chapter on Skewness). In cases of moderately skew (or moderately asymmetrical) distributions the value of the mean, median and mode have the following empirical relationship which is given by Karl Pearson : (i) Mode = Mean—3 (Mean—Median) = Mean—3 Mean + 3 Median = 3 Median — 2 Mean Thus the value of Mode — 3 Median—2 Mean. From this relationship, Mode = Mean—3 (Mean—Median) we can derive (ii) Mean—Mode = 3 (Mean—Median) or (iii) Mean—Median = 1/3 (Mean—Mode) Thus we see that the difference between Mean and Mode is three times, the difference between mean and median. In other words median is closer to the mean than mode. This relationship between Mean (X), Median (M) and Mode (Z) is shown in the following diagram.

OpenStudy (anonymous):

read the link ... interesting !

OpenStudy (dumbcow):

ahh just did...thanks yes formula only works in special case where mean does not equal median and distribution is skewed

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

lucky us ... no idea how else it could be done

OpenStudy (callisto):

I understand dumbcow / Mani 's solution, And the link provided by mathg8 is ... interesting...

OpenStudy (anonymous):

I guess you can take each expression and make it median ... then check the answers .

OpenStudy (callisto):

That formula works in this case

OpenStudy (anonymous):

Mani's solution was nice ...

OpenStudy (callisto):

But the way he derived this formula is... unknown :S

OpenStudy (callisto):

he = Karl Pearson

OpenStudy (mani_jha):

Thanks mathg8!

OpenStudy (anonymous):

@Mani_Jha welcome

OpenStudy (callisto):

Thanks all :)

OpenStudy (anonymous):

np